OpenStudy (anonymous):

write the function f(x)= 2x^2+8x+7 in the form of f(x)= a(x-h)^2 + k and find the vertex

OpenStudy (anonymous):

first complete the square... are u familiar in completing the square..?

OpenStudy (anonymous):

no im not

OpenStudy (anonymous):

i'm not quite sure how to complete the square

OpenStudy (anonymous):

separate 2x^2 + 8x from 7 when u complete the square... u get the half of 8 and square it, then add it to both sides. \[2x^2+8x+(\frac{ 8 }{ 2 })^2 = -7\] \[2x^2+8x+(4)^2 = -7+(4)^2\]

OpenStudy (anonymous):

whats the answer..?

OpenStudy (anonymous):

um.. 2x^2+8x+16=9 ?

OpenStudy (anonymous):

its -9..

OpenStudy (anonymous):

so now factor that..can u do it ..?

OpenStudy (anonymous):

im still confused.

OpenStudy (anonymous):

why is it negative 9?

OpenStudy (anonymous):

ah sorry... its positive... anyway can u factor it in the form a(x-h)^2 + k..?

OpenStudy (anonymous):

and then how do i know what point the vertex would be?

OpenStudy (anonymous):

so the form that it should be is f(x)= 2(x+2)^2 -9 ?